A patient order requires 1.9 L of a 0.5 % benzalkonium chloride antiseptic. You have a 17 % stock solution. How much stock solution will you need to prepare the order?
Answer #1

Let of volume of stock solution required be v ml

1.9L=1900 ml

% of stock solution * volume required = % of solution to be made * volume to be made

17 * v = 0.5 * 1900

v = (0.5*1900)/17

v = 55.88 ml

thus , 55.88 ml of stock solution must be dissolved to 1.9L to obtain 0.5% solution.

If you have any query regarding this you may ask in comments.